The set of all the points (x, y,z) satisfying the equation x−z = z−y represents a

line. Check whether true or false.


1
Expert's answer
2020-02-26T11:49:21-0500

"x-z=z-y\\
x+y-2z=0"

The canonical equation of the plane is ax+by+cz+d=0.

This means that the equation represents a plane, not a line.

Answer: false.